### v1

@@ Abstract
This article discusses strong nuggets (SNs) which means strong interaction condensed matter clusters with a mass of about $10^{6}\,$g. They may originate from the early universe, supernova, pulsar merger event, and so on. Depending on the equation of state, the SNs could be stable and even be one of the candidates for dark matter. In order to detect SNs which hitting the Earth or the Moon at a non-relativistic velocity, a new messenger, the acoustic array, is analysed. The results of the calculations show that the impact signal of an SN can be detected at a distance of about 30 kilometers from the nugget’s trajectory. By using microphone boxes, hydrophones or seismographs to construct an array in the bedrock, ocean or on the Moon, it is possible to reconstruct the velocity, mass, and interacting cross section of SNs, and then constrain also the nature of supra-nuclear matter. The acoustic array can also be used for distributed acoustic sensing of meteorites or earthquakes. The sonar localisation system on the proposed High-energy Underwater Neutrino Telescope (HUNT) is suggested as a pathfinder for acoustic array detection.
@@ 1 Introduction
We are studying the interaction of SNs travelling through the Earth or the Moon at a non-relativistic velocity, and trying to detect the signals with an acoustic array. We perform our calculation using spontaneously magnetized SnNs (Lai & Xu 2016) as an example, but the results could also be meaningful for other types of SNs if the cross sections are adjusted. The physical properties such as radius, mass, number density, and cross section of SNs are introduced in Section 2. Section 3 is relevant to the acoustic detection of an SN, including the penetration depth, the characteristics of the sound waveform generated, and the decay and critical detectable distance of acoustic signals in water and rock media. Discussion and conclusion are presented in Section 4 and Section 5. Unless otherwise stated, we formulate in the cgs Gaussian unit system.
@@ 2 Physical properties of strong nuggets
There are some similar properties among different kinds of SNs. The density of SN is about 1.5 times the density of nuclear matter (Yuan et al. 2024), $\rho_{\rm{SN}}\simeq 4\times 10^{14}\,\rm{g\cdot cm^{-3}}$, with a uniform density at pressure free. We will use $A=10^{30}$ as a typical baryon number of the SN, and can then calculate the mass of the SN, $m_{\rm{SN}}=1.7\times 10^{6}\,(A/10^{30})\,\rm g$, as well as the geometric radius,
$r_{0}=(\dfrac{3m_{\rm{SN}}}{4\pi\rho_{\rm{SN}}})^{1/3}=1.6\times 10^{-3}\,\rm{cm}\cdot(\dfrac{\it{A}}{10^{30}})^{1/3}(\dfrac{\rho_{\rm{SN}}}{10^{14}\,\rm{g\cdot cm^{-3}}})^{-1/3}.$ (1)
Another important parameter is the charge-mass ratio of the nugget, $\delta_{\rm e}=N_{\rm e}/A$, with $N_{\rm e}$ the number of electrons in a single SN. For strange quark matter, we have $\delta_{\rm e}\sim 10^{-7}$ to $10^{-4}$ in the bag model, depending on the strange quark mass $m_{\rm s}$ ($\sim 100$ MeV) and the coupling constant $\alpha_{\rm c}$ (Farhi & Jaffe 1984). For flavour-asymmetric $ud$QM in bulk, one could also have a relatively small $\delta_{\rm e}$ to allow for Landau’s neutralization. For SnM, the ratio $\delta_{\rm e}$ could be of order of $10^{-7}-10^{-4}$, since it would be considered as linked bags of strangeons (Miao et al. 2022). Thanks to this extremely small $\delta_{\rm e}$, SNs can be candidates for dark matter. In this picture, most SNs are virially bound in the Milky Way, moving at a typical velocity of $v=220\,{\rm km}\cdot{\rm s}^{-1}$ relative to Earth in all directions. A small part of the SNs are generated in the supernova or compact star merger process, they may have speed up to order of $v=0.01c$ (Siraj & Loeb 2020). The number density of SN is $n_{\mathrm{SN}}=\rho_{\mathrm{DM}}/m_{\mathrm{SN}}=3.2\times 10^{-31}\,\rm{cm}^{-3}\cdot(\it{A}/\rm{10^{30}})^{-1}$. Then the event rate of SN impact is
$\varphi_{\mathrm{SN}}=n_{\mathrm{SN}}v=7.0\times 10^{-24}\,\rm{s^{-1}cm^{-2}}\cdot(\dfrac{\it{A}}{10^{30}})^{-1}\cdot\dfrac{\it{v}}{220\,\rm{km\cdot{s^{-1}}}}.$ (2)
@@
$\sigma_{\mathrm{SN\_self}}/m_{\mathrm{SN}}=\pi\lambda_{\rm D}^{2}/m_{\mathrm{SN}}=0.9\,\rm{cm}^{2}\cdot(\dfrac{\it{A}}{10^{30}})^{-1},$ (3)
which is consistent with the observational limits of self-interacting dark matter (Carlson et al. 1992; Popolo et al. 2020; Tseng & Yeh 2024; Nadler et al. 2024).
@@
$\sigma_{\mathrm{SN\_geo}}=\pi r_{0}^{2}=8.0\times 10^{-6}\,\rm{cm}^{2}\cdot(\dfrac{\it{A}}{10^{30}})^{2/3}(\dfrac{\rho_{\rm{SN}}}{10^{14}\,\rm{g\cdot cm^{-3}}})^{-2/3},$ (4)
@@
We are concerned about the cross section of SN interacting with ordinary matter below. It is worth noting that an SnN, as an example, could be spontaneously magnetized (Lai & Xu 2016), and we calculate the magnetic momentum of SnN first. The calculation could also be applied to other types of SN if one fixes up the cross section.
An SnN interacts with ordinary matter mainly through its dipole magnetic field, which is generated by the spontaneous magnetization of electrons. We assume that the magnetic moment of strangeons (i.e. quarks) are arranged randomly and cancel each other out, so the contribution to the total magnetic field can be neglected. SnN’s magnetic moment resulting from electric current is neglected due to an extremely small magnetic Reynolds number [^4] However, the dynamo action in massive strangeon stars could be effective, and strangeon magnetars could then form because of the persistent electric current (Xu 2024). , so we will consider only the spin magnetic moment of electrons. As an identical fermion system, the wave function of electrons must be exchange-anti-symmetrical. For example, one can consider the two-electron wave function without spin-orbit coupling, $\Psi(r_{1},r_{2},s_{1},s_{2})=\psi(r_{1},r_{2})\chi(s_{1})\chi(s_{2})$, in which $r_{1}$ and $r_{2}$ are the space coordinates of electrons marked 1 and 2, $s_{1}$ and $s_{2}$ are the spins of electrons. If the total spin is $S=|\vec{s}_{1}+\vec{s}_{2}|=1$, then $\chi(s_{1})\chi(s_{2})$ is exchange-symmetrical, so $\psi(r_{1},r_{2})$ must be exchange-anti-symmetrical. In this case, $\psi(r_{1},r_{2})=0$ when $r_{1}=r_{2}$, the probability that two electrons are in the same position tends to zero. If the total spin is $S=0$ instead, $\psi(r_{1},r_{2})$ is exchange-symmetrical and need not to be zero when $r_{1}=r_{2}$. Because of the Coulomb repulsion between electrons, the interaction energy in $S=1$ case is then lower than that in $S=0$. As a result, spontaneous magnetisation occurs in the electron system inside the SnN, reversing the spins of many electrons, giving a large total spin $S$ and leading to a large magnetic moment $\mu$.
where $\alpha_{\rm em}\equiv e^{2}/(\hbar c)\simeq 1/137$. As $\Delta p\ll p_{\rm F}$, there is only a small fraction of electrons whose spins are parallel when the Coulomb interaction is included for an electron gas. Therefore, the total magnetic momentum reads,
where $\mu_{\rm B}=e\hbar/(2m_{\rm e})$ is the Bohr magneton and $N_{\rm e}$ is the total number of electrons in an SN. The magnetic field on the magnetic equator of the SnN is thus,
$B_{0}=\dfrac{\xi\mu}{r_{0}^{3}}=8.4\times 10^{11}\,\rm{G}\cdot\dfrac{\delta_{\rm e}}{10^{-5}}\cdot\dfrac{\rho_{\rm{SN}}}{10^{14}\,\rm{g\cdot cm^{-3}}},$ (9)
which is close to the typical magnetic fields on pulsars’ surface.
To calculate the cross section of SN,let us see what happens when an SN impacts the ordinary matter, such as soil, water or rock. The SN makes a deposition of its kinetic energy into the surrounding medium. Because its geometric radius ($\sim 10^{-3}\,$cm, see Eq. (1)) could be much larger than the typical distance between molecules, we may treat the medium as continuous. The way SN loses energy is by compressing the medium and transferring momentum to the molecules of the medium. Some of the energy is converted to heat, creating an ionised environment, and some of the energy creates the shock wave as an acoustic signal. In the aerodynamic heating picture, the temperature rises, $\Delta T=v^{2}/2c_{\rm p}$, where $c_{\rm p}$ is the specific isobaric heat capacity. For solid SiO2, $c_{\rm p}=7.0\times 10^{6}\,{\rm erg\cdot(g\cdot K)^{-1}}$. Then substitute $v=220\,\rm{km\cdot s^{-1}}$, we have $\Delta T=7.0\times 10^{7}\,$, which is high enough to ionise the atoms, causing the medium surrounding the impact trajectory to become plasma.
$r_{\rm m}=(\dfrac{B_{0}^{2}r_{0}^{6}}{4\pi\rho_{\rm med}v^{2}})^{1/6}=0.035\,{\rm cm}\cdot(\dfrac{\it{A}}{\mathrm{10^{30}}}\dfrac{\delta_{\rm e}}{\rm{10^{-5}}})^{1/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1/6}(\dfrac{\it{v}}{\rm{220\,km\cdot s^{-1}}})^{-1/3}.$ (10)
It implies that $r_{\rm m}\gg r_{0}$, and $r_{\rm m}/r_{0}=22\cdot(\dfrac{\delta_{\rm e}}{\rm{10^{-5}}}\dfrac{\rho_{\rm{SN}}}{10^{14}\,\rm{g\cdot cm^{-3}}})^{1/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{1/6}(\dfrac{\it{v}}{\rm{220\,km\cdot s^{-1}}})^{-1/3}$ is independent of the baryon number $A$, i.e. the size of SnN. Thanks to the large $r_{\rm m}$, acoustic detection of the SN is more feasible, which will be the subject of Section 3.
The cross section of the magnetosphere of the SnN reads,
$\sigma_{\rm m}=\pi r_{\rm m}^{2}=3.8\times 10^{-3}\,{\rm cm}^{2}\cdot(\dfrac{\it{A}}{\rm{10^{30}}}\dfrac{\delta_{\rm e}}{\rm{10^{-5}}})^{2/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1/3}(\dfrac{\it{v}}{\rm{220\,km\cdot s^{-1}}})^{-2/3},$ (11)
which could be regarded as the maximum effective cross section of the SnN in the ordinary matter medium.
@@
For different kinds of SNs, because of possible absence of the magnetic field, different type of interactions (some kinds of SN may interact with ordinary matter through nuclear force, some through electromagnetic force), and the incompletely momentum transfer from SN to medium particles, the effective cross section could be different. For convenience, we may define the cross section in general
$\sigma_{\rm SN}=k\cdot{(\dfrac{\it{A}}{\rm{10^{30}}})}^{2/3}\cdot 3.8\times 10^{-3}\,{\rm cm}^{2},$ (12)
@@
where the factor $k$ measures the real cross sections of different SNs. Although the relations of $\sigma_{\rm SN}-A$ change in different SN models, we can still scale the cross section with $A^{2/3}$, which is proportional to the geometric section of SN, and roughly set $k$ independent of $A$. Certainly, the cross section $\sigma_{\rm{SN}}$ depends also on the velocity $v$, but for the sake of simplify, we treat $\sigma_{\rm{SN}}$ as a constant when $v$ changes. This is due to the fact that the penetration depth of SN, as described in Section 3, is greater than 10 km. For an acoustic array whose depth is a few kilometres at most, the velocity $v$ of SN does not change significantly.
@@
The range of $k$-value is discussed as following. The cross section of SnN with magnetosphere is assumed to be the largest in all SN models, indicating the upper limit of $k$-value. In this case, the factor about $\rho_{\rm med}$ and $v$ is order 1, so the maximum $k$-value is when $\delta_{\rm e}$ reaches the largest, $10^{-4}$, and the magnetosphere transfers momentum completely, leading to that of $k=10^{2/3}$. The geometric section could be much smaller, and the incompletely momentum transfer would aditionally further reduce the effective cross section. In this case, $\sigma_{\rm SN}\leq\pi r_{0}^{2}$, $k\leq 1/22^{2}=2.1\times 10^{-3}$. Therefore, we take $10^{-3}<k<10^{1}$ in the calculations below.
@@ 3 Acoustic detection
Because the velocity of SN is non-relativistic, the shower effect may not be significant, but the mechanical shock wave, i.e. acoustic wave, may be detectable.
First, we calculate the penetration depth of an SN. The energy deposition in unit penetration depth is $-{\mathrm{d}E_{\rm k}}/{\mathrm{d}x}=F_{\rm{drag}}=\sigma_{\rm{SN}}\cdot\rho_{\mathrm{med}}v^{2}/2$, with $E_{\rm k}=m_{\mathrm{SN}}v^{2}/2$. Solving this ordinary differential equation, we can obtain $x=8.8\times 10^{8}\,{\rm cm}\cdot k^{-1}(\dfrac{\it{A}}{\rm{10^{30}}})^{1/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1}\ln(\dfrac{220\,{\rm km\cdot s^{-1}}}{v(x)})$. It is evident that the velocity decays exponentially with penetration depth, and the characteristic decay depth is thus
$x_{\rm{pene}}=8.8\times 10^{8}\,{\rm cm}\cdot k^{-1}(\dfrac{\it{A}}{\rm{10^{30}}})^{1/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1}.$ (13)
@@
When SNs impact the Earth, they will first pass through the atmosphere. To evaluate whether an SN can reach the ground, we make the approximation that the atmosphere has uniform density $\rho_{\rm med}=1.2\times 10^{-3}\,\rm{g\cdot cm^{-3}}$ and has a thickness of $x_{\rm atm}=k_{\rm B}T/\bar{m}_{\rm atm}g=8.8\,$km. Let $x_{\rm{pene}}>x_{\rm atm}$, one can derive easily that $A>k^{3}\cdot 1.7\times 10^{12}$. All the SNs with $10^{25}<A<10^{35}$ and $10^{-3}<k<10^{1}$ can reach the ground, we ignore the influence of the Earth’s atmosphere below.
@@
$\begin{split}t_{\rm s}&=1.82\cdot\frac{2\sqrt{\sigma_{\rm SN}}}{c_{\rm s}}(\dfrac{vR}{c_{\rm s}\cdot\sqrt{\sigma_{\rm SN}}})^{1/4}\\ &=0.31\,{\mathrm{ms}}\cdot k^{3/8}(\dfrac{A}{10^{30}}\dfrac{R}{1\,{\mathrm{km}}}\dfrac{v}{220\,{\mathrm{km\cdot s^{-1}}}})^{1/4}(\dfrac{c_{\rm s}}{1\,{\mathrm{km\cdot s^{-1}}}})^{-5/4},\end{split}$ (14)
where $\sqrt{\sigma_{\rm SN}}$ is the effective size of the SN “bullet” in the ordinary medium, $R$ is the distance from the SN trajectory, $c_{\rm s}$ is the sound speed in the medium. The frequency of shock wave is order of $f_{\rm s}=1/t_{\rm s}$. For $10^{25}<A<10^{35}$, $10^{-3}<k<10^{1}$, $R=1\,$km, and $c_{\rm s}=3\,{\rm km\cdot s^{-1}}$, we have $f_{\rm s}$ from $\sim 0.30$ kHz to $\sim 3.0$ MHz. SNs with larger ${\sigma_{\rm SN}}$ could have lower frequency, but the environment noise is stronger, and the event rate is lower. SNs with smaller ${\sigma_{\rm SN}}$ have higher frequency, but this high frequency sound wave would decay fast in the medium. When the distance $R$ increases, the frequency decreases as $R^{-1/4}$. Taking all these factors into account, we can use civilian microphones with a sampling rate of 192 kHz to detect the SN signal whose frequency is lower than 96 kHz (half of 192 kHz). We could also use existing earthquake monitoring stations, but the sampling rate of existing seismographs, 1 Hz to 1 kHz, is relatively low (Tian & Liang 2013; Goldman et al. 2020; Chen et al. 2022; Darzi et al. 2024). It is necessary to develop advanced monitors to capture the SN signals at high frequency.
@@
in which $\mathrm{d}E_{\rm s0}$ is the acoustic energy per unit penetration depth at $R\rightarrow 0$. The value of $R_{0}$ varies with different media, temperature, and wave frequency. For example, we have $R_{0}=31\,$km for a 2 kHz sound wave in $5\,^{\circ}$ ocean water, while $R_{0}=1.1\,$km for 20 kHz (Waite 2002). The propagation of pressure waves in the solid crust of the Earth or the Moon is similar to that of seismic body waves. The uniformity of the crust is less than that of the ocean water, and the compression wave and the shear wave will transfer each other at crust-air or crust-mantle interfaces during propagation, so even though the crust is solid, $R_{0}$ will not get much larger. If the seismic wave attenuation data in Preliminary Reference Earth Model (Dziewonski & Anderson 1981) is used in calculation, for a 2 kHz compression wave, $R_{0}=30$ km; for a 2 kHz shear wave, $R_{0}=0.19$ km. For estimation, we take the value of $R_{0}=1$ km. The sound wave will travel to about several times of $R_{0}$, and then the noise will dominate.
We then do some more detailed calculations, taking into account the changes of $A$ and $k$. By equating the pressures of sound wave and of environment noise $p_{\rm s}=p_{\rm noise}$, one may estimate the critical detectable distance $R_{\rm crit}$. We need equations $-{\mathrm{d}E_{\rm k}}/{\mathrm{d}x}=\sigma_{\rm{SN}}\cdot\rho_{\mathrm{med}}v^{2}/2$, $\mathrm{d}E_{\rm s}=\mathrm{d}E_{\rm s0}\exp(-R/R_{0})=(-\mathrm{d}E_{\rm k})\exp(-R/R_{0})$, Equation (15), and Equation (14). The pressure of background noise includes human and animal activity, tiny quakes, and weather. At the Salt Lake, VanDevender et al. (2017) measured $p_{\rm noise}\simeq 1\,{\rm erg}\cdot{\rm cm}^{-3}$, about 74 dB (sound pressure level referenced to 20 µPa). We use this value to represent the background noise level around the acoustic array. The critical detectable distance $R_{\rm crit}$ satisfies,
$\begin{split}\exp(-\dfrac{R_{\rm crit}}{1\,\mathrm{km}})&=\dfrac{4\pi R_{\rm crit}p_{\rm noise}^{2}t_{\rm s}}{\rho_{\mathrm{med}}^{2}c_{\rm s}\sigma_{\rm SN}v^{2}}\\ &=\exp(-19.88)\cdot(\dfrac{A}{10^{30}})^{-5/12}(\dfrac{R_{\rm crit}}{1\,\mathrm{km}})^{5/4}(\dfrac{p_{\rm noise}}{1\,\rm{erg\cdot cm^{-3}}})^{2}(\dfrac{v}{220\,\rm{km\cdot s^{-1}}})^{-7/4}\\ &\cdot k^{-5/8}(\dfrac{c_{\rm s}}{1\,\rm{km\cdot s^{-1}}})^{-9/4}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-2}.\end{split}$ (17)

### v2

@@ Abstract
This article discusses strong nuggets (SNs) which means strong interaction condensed matter with a mass of about $10^{6}\,$g. They may originate from the early universe, supernova, pulsar merger event, and so on. Depending on the equation of state, the SNs could be stable and even be one of the candidates for dark matter. In order to detect SNs which hitting the Earth or the Moon at a non-relativistic velocity, a new messenger, the acoustic array, is analysed. The results of the calculations show that the impact signal of an SN can be detected at a distance of about 30 kilometers from the nugget’s trajectory. By using microphone boxes, hydrophones or seismographs to construct an array in the bedrock, ocean or on the Moon, it is possible to reconstruct the velocity, mass, and interacting cross section of SNs, and then constrain also the nature of supra-nuclear matter. The acoustic array can also be used for distributed acoustic sensing of meteorites or earthquakes. The sonar localisation system on the proposed High-energy Underwater Neutrino Telescope (HUNT) is suggested as a pathfinder for acoustic array detection.
@@ 1 Introduction
We are studying the interaction of SNs travelling through the Earth or the Moon at a non-relativistic velocity, and trying to detect the signals with an acoustic array. We perform our calculation using spontaneously magnetized SnNs (Lai & Xu 2016) as an example, but the results could also be meaningful for other kinds of SNs if the cross sections are adjusted. The physical properties such as radius, mass, number density, and cross section of SNs are introduced in Section 2. Section 3 is relevant to the acoustic detection of an SN, including the penetration depth, the characteristics of the sound waveform generated, and the decay and critical detectable distance of acoustic signals in water and rock media. Discussion and conclusion are presented in Section 4 and Section 5. Unless otherwise stated, we formulate in the cgs Gaussian unit system.
@@ 2 Physical properties of strong nuggets
There are some similar properties among different kinds of SNs. The density of SN is about 1.5 times the density of nuclear matter (Yuan et al. 2025), $\rho_{\rm{SN}}\simeq 4\times 10^{14}\,\rm{g\cdot cm^{-3}}$, with a uniform density at pressure free. We will use $A=10^{30}$ as a typical baryon number of the SN, and can then calculate the mass of the SN, $m_{\rm{SN}}=1.7\times 10^{6}\,\rm g\cdot(A/10^{30})$, as well as the geometric radius,
$\begin{split}&r_{0}=(\dfrac{3m_{\rm{SN}}}{4\pi\rho_{\rm{SN}}})^{1/3}\\ &=1.6\times 10^{-3}\,\rm{cm}\cdot(\dfrac{\it{A}}{10^{30}})^{1/3}(\dfrac{\rho_{\rm{SN}}}{10^{14}\,\rm{g\cdot cm^{-3}}})^{-1/3}.\end{split}$ (1)
Another important parameter is the charge-mass ratio of the nugget, $\delta_{\rm e}=N_{\rm e}/A$, with $N_{\rm e}$ the number of electrons in a single SN. For strange quark matter, we can calculate that $\delta_{\rm e}\sim 10^{-6}$ to $10^{-5}$ in the MIT bag model, depending on the strange quark mass $m_{\rm s}$ (taken as $\sim 95$ MeV) and the coupling constant $\alpha_{\rm c}$ (Farhi & Jaffe 1984). For flavour-asymmetric $ud$QM in bulk, one could also have a relatively small $\delta_{\rm e}$ to allow for Landau’s neutralization. For SnM, the ratio $\delta_{\rm e}$ could be of order of $10^{-6}-10^{-5}$, since it would be considered as linked bags of strangeons (Miao et al. 2022). Thanks to this extremely small $\delta_{\rm e}$, SNs can be candidates for dark matter. In this picture, most SNs in the vicinity of the solar system are virially bound to the Milky Way, moving at a typical velocity of $v=220\,{\rm km}\cdot{\rm s}^{-1}$ relative to Earth in all directions. A small part of the SNs are generated in the supernova or compact star merger process, they may have speed up to order of $v=0.01c$ (Siraj & Loeb 2020). The number density of SN is $n_{\mathrm{SN}}=\rho_{\mathrm{DM}}/m_{\mathrm{SN}}=3.2\times 10^{-31}\,\rm{cm}^{-3}\cdot(\it{A}/\rm{10^{30}})^{-1}$. Then the event rate of SN impact is
$\begin{split}&\varphi_{\mathrm{SN}}=n_{\mathrm{SN}}v\\ &=7.0\times 10^{-24}\,\rm{s^{-1}cm^{-2}}\cdot(\dfrac{\it{A}}{10^{30}})^{-1}\cdot\dfrac{\it{v}}{220\,\rm{km\cdot{s^{-1}}}}.\end{split}$ (2)
@@
$\begin{split}&\sigma_{\mathrm{SN\_self}}/m_{\mathrm{SN}}=\pi\lambda_{\rm D}^{2}/m_{\mathrm{SN}}\\ &=0.9\,\rm{cm}^{2}\cdot\rm{g}^{-1}\cdot(\dfrac{\it{A}}{10^{30}})^{-1},\end{split}$ (3)
which is consistent with the observational limits of self-interacting dark matter (Carlson et al. 1992; Popolo et al. 2020; Tseng & Yeh 2024; Nadler et al. 2025).
@@
$\begin{split}&\sigma_{\mathrm{SN\_geo}}=\pi r_{0}^{2}\\ &=8.0\times 10^{-6}\,\rm{cm}^{2}\cdot(\dfrac{\it{A}}{10^{30}})^{2/3}(\dfrac{\rho_{\rm{SN}}}{10^{14}\,\rm{g\cdot cm^{-3}}})^{-2/3},\end{split}$ (4)
@@
For different kinds of SNs, because of the possible magnetic field, different types of interactions (some kinds of SN may interact with ordinary matter through nuclear force, some through electromagnetic force), and the incompletely momentum transfer from SN to medium particles, the effective cross section could be different. See Appendix A for detailed calculations. For convenience, we may define the cross section in general
$\sigma_{\rm SN}=k\cdot{(\dfrac{\it{A}}{\rm{10^{30}}})}^{2/3}\cdot 8.0\times 10^{-6}\,{\rm cm}^{2}.$ (5)
@@
Here, the factor $k$ measures the real cross sections relative to the geometric cross sections of different SNs. Although the relations of $\sigma_{\rm SN}-A$ change in different SN models, we can still scale the cross section with $A^{2/3}$, which is proportional to the geometric section of SN, and roughly set $k$ independent of $A$. Certainly, the cross section $\sigma_{\rm{SN}}$ depends also on the velocity $v$, but for the sake of simplify, we treat $\sigma_{\rm{SN}}$ as a constant when $v$ changes. This is due to the fact that the penetration depth of SN, as described in Section 3, is greater than 10 km. For an acoustic array whose depth is a few kilometres at most, the velocity $v$ of SN does not change significantly.
@@
The range of $k$-value is discussed as following. On the one hand, the geometric cross section is a benchmark for the value of $k$. For an SN that interacts directly with ordinary matter via its geometric cross section, $k=1$. The incompletely momentum transfer would reduce the effective cross section, but not by orders of magnitude, so the lower limit of $k$ is taken to be $10^{-1}$. On the other hand, the cross section of the $ud$-quark nugget with magnetosphere is assumed to be the largest in all SN models, indicating the upper limit of $k$-value, $k=10^{3}$. Therefore, we take $10^{-1}<k<10^{3}$ in the calculations below. Table 1 shows the $k$ values for different kinds of SNs.
Table 1: The $k$ Value for Different Kinds of SNs.
Type of the SN $k$
strange quark/antiquark nugget $10^{1}\sim 10^{2}$
$ud$-quark/antiquark nugget $\sim 10^{3}$
axion-quark/antiquark nugget $0.1\sim 10^{3}$
strangeon nugget $10^{1}\sim 10^{2}$
@@ 3 Acoustic detection
As SNs travel at non-relativistic velocities when they impact the Earth, the shower effect may not be significant. However, the mechanical shock wave (i.e. the acoustic wave) may be detectable. Due to the geometric radius of an SN (about $10^{-5}\sim 10^{-1}\,$cm, see Eq. 1) is much larger than the typical distance between molecules (about $10^{-8}\sim 10^{-7}\,$cm), we can treat the medium as continuous, the hydrodynamic equations can be applied.
First, we calculate the penetration depth of an SN. The energy deposition in unit penetration depth is $-{\mathrm{d}E_{\rm k}}/{\mathrm{d}x}=F_{\rm{drag}}=\sigma_{\rm{SN}}\cdot\rho_{\mathrm{med}}v^{2}/2$, with $E_{\rm k}=m_{\mathrm{SN}}v^{2}/2$. Solving this ordinary differential equation, we can obtain $x=4.2\times 10^{11}\,{\rm cm}\cdot k^{-1}(\dfrac{\it{A}}{\rm{10^{30}}})^{1/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1}\ln(\dfrac{220\,{\rm km\cdot s^{-1}}}{v(x)})$. It is evident that the velocity decays exponentially with penetration depth, and the characteristic decay depth is thus
$\begin{split}x_{\rm{pene}}&=4.2\times 10^{11}\,{\rm cm}\\ &\cdot k^{-1}(\dfrac{\it{A}}{\rm{10^{30}}})^{1/3}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-1}.\end{split}$ (6)
@@
When SNs impact the Earth, they will first pass through the atmosphere. To evaluate whether an SN can reach the ground, we make the approximation that the atmosphere has uniform density $\rho_{\rm med}=1.2\times 10^{-3}\,\rm{g\cdot cm^{-3}}$ and has a thickness of $x_{\rm atm}=k_{\rm B}T/\bar{m}_{\rm atm}g=8.8\,$km. Let $x_{\rm{pene}}>x_{\rm atm}$, one can derive easily that $A>k^{3}\cdot 1.6\times 10^{4}$. All the SNs with $10^{25}<A<10^{35}$ and $10^{-1}<k<10^{3}$ can reach the ground, we ignore the influence of the Earth’s atmosphere below.
@@
$\begin{split}t_{\rm s}&=1.82\cdot\frac{2\sqrt{\sigma_{\rm SN}}}{c_{\rm s}}(\dfrac{vR}{c_{\rm s}\cdot\sqrt{\sigma_{\rm SN}}})^{1/4}\\ &=0.031\,{\mathrm{ms}}\cdot k^{3/8}(\dfrac{A}{10^{30}}\dfrac{R}{1\,{\mathrm{km}}}\dfrac{v}{220\,{\mathrm{km\cdot s^{-1}}}})^{1/4}\\ &\cdot(\dfrac{c_{\rm s}}{1\,{\mathrm{km\cdot s^{-1}}}})^{-5/4},\end{split}$ (7)
where $\sqrt{\sigma_{\rm SN}}$ is the effective size of the SN “bullet” in the ordinary medium, $R$ is the distance from the SN trajectory, $c_{\rm s}$ is the sound speed in the medium. The frequency of shock wave is order of $f_{\rm s}=1/t_{\rm s}$. For $10^{25}<A<10^{35}$, $10^{-1}<k<10^{3}$, $R=1\,$km, and $c_{\rm s}=3\,{\rm km\cdot s^{-1}}$, we have $f_{\rm s}$ from $\sim 0.54$ kHz to $\sim 5.4$ MHz.
[…]
Signals from SNs with larger ${\sigma_{\rm SN}}$ could have lower frequency, but the environment noise is stronger in the low-frequency region, and the event rate of a larger ${\sigma_{\rm SN}}$ is lower. Signals from SNs with smaller ${\sigma_{\rm SN}}$ have higher frequency, but this high frequency sound wave would decay fast in the medium. When the distance $R$ increases, the frequency decreases. Taking all these factors into account, we can use civilian microphones with a sampling rate of 192 kHz to detect the SN signal whose frequency is lower than 96 kHz (half of 192 kHz). We could also use existing earthquake monitoring stations, but the sampling rate of existing seismographs, 1 Hz to 1 kHz, is relatively low (Tian & Liang 2013; Goldman et al. 2020; Chen et al. 2022; Darzi et al. 2024). It is necessary to develop advanced monitors to capture the SN signals at high frequency.
@@
in which $\mathrm{d}E_{\rm s0}$ is the acoustic energy per unit penetration depth at $R\rightarrow 0$. The value of $R_{0}$ varies with different media, temperature, and wave frequency. For example, we have $R_{0}=31\,$km for a 2 kHz sound wave in $5\,^{\circ}$ ocean water, while $R_{0}=1.1\,$km for 20 kHz (Waite 2002). The propagation of pressure waves in the solid crust of the Earth or the Moon is similar to that of seismic body waves. The uniformity of the crust is less than that of the ocean water, and the compression wave and the shear wave will transfer each other at crust-air or crust-mantle interfaces during propagation, so even though the crust is solid, $R_{0}$ will not get much larger. If the seismic wave attenuation data in Preliminary Reference Earth Model (Dziewonski & Anderson 1981) is used in calculation, for a 2 kHz compression wave, $R_{0}=30$ km; for a 2 kHz shear wave, $R_{0}=0.19$ km. For the purpose of estimation and averaging, we take the value of $R_{0}=1$ km. The sound wave will travel to about several times of $R_{0}$, and then the noise will dominate.
We then do some more detailed calculations, taking into account the changes of $A$ and $k$. By equating the pressures of sound wave and of environment noise $p_{\rm s}=p_{\rm noise}$, one may estimate the critical detectable distance $R_{\rm crit}$. We need the drag equation due to ram pressure $-{\mathrm{d}E_{\rm k}}/{\mathrm{d}x}=\sigma_{\rm{SN}}\cdot\rho_{\mathrm{med}}v^{2}/2$, the conversion of kinetic energy into acoustic energy and decay of acoustic energy $\mathrm{d}E_{\rm s}=\mathrm{d}E_{\rm s0}\exp(-R/R_{0})=(-\mathrm{d}E_{\rm k})\exp(-R/R_{0})$, the duration Equation (7) of the sound pressure, and Equation (8) for calculating sound energy from integrating the sound pressure. The pressure of background noise includes human and animal activity, tiny quakes, and weather. At the Salt Lake, VanDevender et al. (2017) measured $p_{\rm noise}\simeq 1\,{\rm erg}\cdot{\rm cm}^{-3}$, about 74 dB (sound pressure level referenced to 20 µPa). We use this value to represent the background noise level around the acoustic array. The critical detectable distance $R_{\rm crit}$ satisfies,
$\begin{split}\exp(-\dfrac{R_{\rm crit}}{1\,\mathrm{km}})&=\dfrac{4\pi R_{\rm crit}p_{\rm noise}^{2}t_{\rm s}}{\rho_{\mathrm{med}}^{2}c_{\rm s}\sigma_{\rm SN}v^{2}}\\ &=\exp(-16.03)\cdot(\dfrac{A}{10^{30}})^{-5/12}(\dfrac{R_{\rm crit}}{1\,\mathrm{km}})^{5/4}\\ &\cdot(\dfrac{p_{\rm noise}}{1\,\rm{erg\cdot cm^{-3}}})^{2}(\dfrac{v}{220\,\rm{km\cdot s^{-1}}})^{-7/4}\\ &\cdot k^{-5/8}(\dfrac{c_{\rm s}}{1\,\rm{km\cdot s^{-1}}})^{-9/4}(\dfrac{\rho_{\rm med}}{1\,\rm{g\cdot cm^{-3}}})^{-2}.\end{split}$ (10)